Admission to mammoth cave is $12 for adults and $8 for children. One day, 575 people entered the cave paying a total of $5,600. How many adults and how many childrenentered the cave?
Answer #1

12x + 8(575-x) =5600

4x = 5600-4600

x =1000/4

x =250 adults
children = 325 children

adults =250
Answer #2

Let x be the adults and y children
12x+8y = 5600
and
x+y =575
Solving these two equations we get,

x= 250 and y= 325
Answer #3
12x + 8(575-x) = 5600
hence x = 250 = adults
children = 325